find the range variance and standard deviation of 2.1,3.0,1.9,1.5,4.8

1
Expert's answer
2022-01-10T16:42:46-0500

Sorted Data Set: "1.5, 1.9, 2.1, 3.0, 4.8"


"Range=4.8-1.5=3.3"

"mean=\mu=\dfrac{1}{5}(1.5+1.9+2.1+3.0+4.8)"

"=2.66"

"Var(x)=\sigma^2=\dfrac{1}{5}((1.5-2.66)^2+(1.9-2.66)^2"

"+(2.1-2.66)^2+(3.0-2.66)^2+(4.8-2.66)^2)"

"=1.3864"

"\sigma=\sqrt{\sigma^2}=\sqrt{1.3864
}\approx1.177455"